OpenStudy (x__hazel__x):

Moira went bowling five times during the summer. Her scores were 195, 190, 190, 250, and 240. If she wanted to misrepresent the data to impress people with how well she bowled during the summer which measure would she most likely use? A. range B. mode C. median D. mean

OpenStudy (x__hazel__x):

@jim_thompson5910 could you please help? :)

OpenStudy (ciarán95):

We want to find which of these will give us the highest score..... Range = Maximum value - Minimum value in the data set = 250 - 190 = 60 Mode = The value which occurs most often in the data set (i.e. the score with the most number of duplicates in the data). Median = The 'middle' value on the data set. In this case, we order the values from smallest to largest and try to pick it out: 190, 190, 195, 240, 250 We have 5 values, so the middle one will be the 3rd value from the left and from the right too. Mean: To calculate the mean, we add up all the values in the data set and divide by the number of values in the set (in this case, 5). \[Mean = \frac{ 190 + 190 + 195 + 240 + 250 }{ 5 }\] I've worked out the range for you, so hopefully you'll be able to finish off finding the median, mode and mean for yourself. Whichever is the largest will give you the correct answer to the question. Hope that helps :)

OpenStudy (x__hazel__x):

thank you! from what you wrote, I think it may be Mean

OpenStudy (ciarán95):

Yes I think you're right @x__hazel__x

OpenStudy (x__hazel__x):

Thank you!